OpenStudy (anonymous):

4m times the square root of m times the square root of 5m

OpenStudy (anonymous):

4m times sqrt 5

OpenStudy (anonymous):

so the answer would be what

OpenStudy (anonymous):

what is the complete equation?

OpenStudy (lyrae):

\[4m \times \sqrt{5m} \times \sqrt{5m} = 4m \times \sqrt{5m}^{2} = 4m \times 5m = 9m\]Because\[\sqrt{a} \times \sqrt{a} = a ^{\frac{ 1 }{ 2 }} \times a ^{\frac{ 1 }{ 2 }} = a ^{\frac{ 1 }{ 2 } + \frac{ 1 }{ 2 }} = a ^{1} = a\]